OpenStudy (anonymous):

What is the sum of the arithmetic sequence 3, 11, 19 ..., if there are 26 terms? A. 1,222 B. 1,261 C. 2,486 D. 2,678

Parth (parthkohli):

\( \color{Black}{\Rightarrow 0.5n(a_1 + a_n) }\) First you find \(a_{26}\)

OpenStudy (anonymous):

what numbers do I plug in where?

Parth (parthkohli):

To find a26: \( \color{Black}{\Rightarrow a_{26} = 3 + 8(26 - 1) }\) Then plug a26 in an

OpenStudy (anonymous):

i don't understand.. a26=225 now what?

Parth (parthkohli):

\( \color{Black}{\Rightarrow a_{26} = 3 + 200 }\) \( \color{Black}{\Rightarrow a_{26} = 203 }\) Back to the formula. \( \color{Black}{\Rightarrow S = 0.5(26)(3 + 203) }\)

OpenStudy (anonymous):

so the answer is D?

Parth (parthkohli):

Yes